OpenStudy (anonymous):

solve -3(2w-1)=5w-19 please show work

jimthompson5910 (jim_thompson5910):

-3(2w-1)=5w-19 -6w+3=5w-19 -6w=5w-19-3 -6w-5w=-19-3 -11w=-19-3 -11w=-22 w=-22/(-11) w=2